# Graph Transformers without Positional Encodings

### v2

@@ 3.3 Eigenformer Architecture
For a layer $l$ of the transformer, input node features $x_{i}^{l}\in\mathbb{R}^{d_{\mathcal{V}}}$ for $i=1,...,|\mathcal{V}|$ and input edge features $e_{ij}\in\mathbb{R}^{d_{\mathcal{E}}}$ for $i,j=1,...,|\mathcal{V}|$, the node feature propagation is performed using:
$\alpha^{l}[i,j]=softmax_{j\in\mathcal{V}}\Bigl{[}\phi_{1}^{l}\bigl{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\phi_{2}^{l}(\lambda_{k})\bigl{)}\Bigl{]}$ (6)
@@
As in Kreuzer et al. [2021], we do not propagate edge features due to the added complexity and marginal benefits in performance. Further, we found degree scaling followed by batch normalization to be highly effective in incorporating node-degree information, as reported in Ma et al. [2023]. The final architecture of a layer of our proposed graph transformer, dubbed Eigenformer is illustrated in Figure 1(b) along with the attention mechanism in Figure 1(a).
Table 1: Test performance on four benchmarks from Dwivedi et al. [2020]. Due to limited training budget, we only show our results for a single training run while other baselines report mean $\pm$ standard deviation numbers of 4 runs with different random seeds. Color codes denote top first, second, third results.
Model ZINC MNIST PATTERN CLUSTER
@@
GCN 0.367$\pm$.011 90.705$\pm$.218 71.892$\pm$.334 68.498$\pm$.976
GIN 0.526$\pm$.051 96.485$\pm$.252 85.387$\pm$.136 64.716$\pm$1.553
GAT 0.384$\pm$.007 95.535$\pm$.205 78.271$\pm$.186 70.587$\pm$.447
GatedGCN 0.282$\pm$.015 97.340$\pm$.143 85.568$\pm$.088 73.840$\pm$.326
GatedGCN-LSPE 0.090$\pm$.001 - - -
PNA 0.188$\pm$.004 97.940$\pm$.12 - -
DGN 0.168$\pm$.003 - 86.680$\pm$.034 -
GSN 0.101$\pm$.010 - - -
CIN 0.079$\pm$.006 - - -
CRaW1 0.085$\pm$.004 97.944$\pm$.050 - -
GIN-AK+ 0.080$\pm$.001 - 86.850$\pm$.057 -
SAN 0.139$\pm$.006 - 86.581$\pm$.037 76.691$\pm$.65
Graphormer 0.122$\pm$.006 - - -
K-Subgraph SAT 0.094$\pm$.008 - 86.848$\pm$.037 77.856$\pm$.104
EGT 0.108$\pm$.009 98.173$\pm$.087 86.821$\pm$.020 79.232$\pm$.348
Graphormer-URPE 0.086$\pm$.007 - - -
Graphormer-GD 0.081$\pm$.009 - - -
GPS 0.070$\pm$.004 98.051$\pm$.126 86.685$\pm$.059 78.016$\pm$.180
GRIT 0.059$\pm$.002 98.108$\pm$.111 87.196$\pm$.076 80.026$\pm$.277
EXPHORMER - 98.55$\pm$.039 86.740$\pm$.015 78.070$\pm$.037
Eigenformer (ours) 0.104 96.178 90.761 76.683
4 Experimental Results
We evaluate our proposed architecture on four benchmarks from the widely used Benchmarking GNNs work in Dwivedi et al. [2020], viz. ZINC, MNIST, PATTERN and CLUSTER. These four datasets include different graph learning objectives like node classification, graph classification and graph regression. They also cover widely different domains ranging from chemistry and computer vision to synthetically generated graphs. Further, we report numbers on two benchmarks from the Long-Range Graph Benchmark (LRGB) [Dwivedi et al., 2022b] to gauge how well our approach does in capturing long-range dependencies. We believe there is room for improvement in the reported results through a more extensive hyperparameter search when more training resources are available. Further details about the experimental setup can be found in appendix B.
Baselines: We compare our method to different SOTA GNN methods including (hybrid) Graph Transformers: GraphGPS Rampášek et al. [2022], GRIT Ma et al. [2023], EGT Hussain et al. [2022], SAN Kreuzer et al. [2021], Graphormer Ying et al. [2021], K-Subgraph SAT Chen et al. [2022b] and EXPHORMER Shirzad et al. [2023], MP-GNNs: GCN Kipf and Welling [2017], GAT Veličković et al. [2018], GIN Xu et al. [2019], PNA Corso et al. [2020], GatedGCN Bresson and Laurent [2018], DGN Beaini et al. [2021], GSN Bouritsas et al. [2023] and other performant models like CIN Bodnar et al. [2021], CRaW1 Tönshoff et al. [2021] and GIN-AK+ Zhao et al. [2021].
Table 2: Test performance on two benchmarks from long-range graph benchmarks (LRGB) [Dwivedi et al., 2022b]. Due to limited training budget, we only show our results for a single training run while other baselines report mean $\pm$ standard deviation numbers of 4 runs with different random seeds. Color codes denote top first, second, third results.
@@ 4 Experimental Results
Eigenformer (ours) 0.6234 0.2749
Figure 3: (Smoothed) similarity $\sigma_{k}$ vs eigenvalue $\lambda_{k}$
@@
(c) PATTERN
(d) CLUSTER
(e) Peptides-func/struct
@@
Figure 3: (Smoothed) similarity $\sigma_{k}$ vs eigenvalue $\lambda_{k}$
@@
Benchmarking GNNs [Dwivedi et al., 2020]: Table 1 compares our results to the baseline methods under the parameter budget used in Rampášek et al. [2022] ($\sim$500K parameters for ZINC, PATTERN and CLUSTER, $\sim$100K parameters for MNIST). Eigenformer routinely beats SOTA MP-GNNs and approaches the performance of popular Graph Transformers that use different PEs like SAN [Kreuzer et al., 2021], Subgraph-SAT [G et al., 2023], Graphormer [Ying et al., 2021] etc. On the PATTERN dataset, we set a new benchmark with the best accuracy among all compared models.
Long-range graph benchmark [Dwivedi et al., 2022b]: To evaluate the ability to learn long-range dependencies, we report numbers on two benchmarks from the LRGB work [Dwivedi et al., 2022b] viz. Peptides-func (10-label graph classification) and Peptides-struct (11-task graph regression) in Table 2. Again, Eigenformer beats all the MP-GNN models easily and is competitive with the SOTA Transformers demonstrating its ability to capture long-range interactions in graph data.
(a) $\phi_{1}$
(b) $\phi_{2}$
(c) Adjacency matrix
(d) Layer 1 attention
(e) Layer 4 attention
(f) Layer 10 attention
Figure 4: $\phi_{1}$, $\phi_{2}$ and attention matrices for an example graph from ZINC dataset
@@
Visualization of node similarities: It is instructive to visualize the difference in the distribution of similarities (defined in 4) between adjacent and non-adjacent nodes. Figure 2(e) shows the (smoothed) average similarity distribution w.r.t frequency for all datasets used in our work. On average, adjacent nodes have higher similarities at the lower end of the spectrum which change inversely with frequency. Non-adjacent nodes on the other hand have smaller similarities at smaller frequencies which rise towards the higher end of the spectrum. We posit that these distributional differences allow the attention mechanism to learn to discern local connections from long-range connections.
Visualization of $\phi_{1}$, $\phi_{2}$ and attention matrices: Figures 3(a) and 3(b) show the learned functions $\phi_{1}^{l}$ and $\phi_{2}^{l}$ for $l=1,...,12$ for the ZINC dataset. The non-linear functions learnt by the model are most likely to ensure that the attention matrix focuses on the required connections/neighborhood as explained in the proof of Proposition 1. For example, Figure 3(d) shows the learned Layer 1 attention matrix for an example graph, which apparently tries to focus more on the 1-hop neighborhood interactions similar to the adjacency matrix shown in Figure 3(c). Figure 3(f) showing Layer 10 attention on the other hand, seems to focus on a larger ($\sim 5$-hop) neighborhood. Finally, as the complex interaction attention matrix in Figure 3(e) shows, $k$-hop neighbor interactions are not the only ones representable by our proposed attention mechanism.
@@ 5 Conclusion
Observing the wide variety of positional encoding schemes for Graph Transformers in the literature, we asked the question whether such encodings are essential to ace graph representational learning objectives. We introduced a novel spectrum-aware attention (SAA) mechanism that incorporates structural graph inductive biases while remaining flexible enough to learn from features in the data. We proposed Eigenformer, a Graph Transformer with SAA at its core, that can be used for a variety of graph learning objectives with different task-specific heads. We empirically test the feasibility of our method through experiments on a number of common GNN benchmarks and find that it performs comparably to many popular GNN techniques using different schems for PE generation, even beating the state-of-the-art on the PATTERN node-classification benchmark.
Several open challenges remain however, including computational complexity of the SAA mechanism with its $O(N^{3})$ space and compute requirements. Recent work on sparse Graph Transformers using expander graphs and sparse attention with virtual/global nodes [Shirzad et al., 2023] along with the use of a subset of frequencies rather than the whole spectrum, may alleviate these issues. We hope this work provides the impetus to the broader graph-representational learning research community to pursue the development of other GNN architectures without the use of PEs.
Appendix A Proof of Proposition 1
@@
$L_{sym}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{sym}$
Since $L_{sym}$ is real and symmetric, it admits an eigendecomposition as follows:
$L_{sym}=U\Lambda U^{-1}$
where the columns of $U=[u_{1},u_{2},...,u_{|\mathcal{V}|}]$ are the right eigenvectors of $L_{sym}$. It follows that,
$A_{sym}=I-U\Lambda U^{-1}=U(I-\Lambda)U^{-1}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}(1-\lambda_{i})$ (11)
@@
$\sum_{k=0}^{m}\theta_{k}A^{k}_{sym}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}\sum_{k=0}^{m}\theta_{k}(1-\lambda_{i})^{k}$ (12)
The continuous functions $\phi_{1}(x)=x$ and $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ in 5 can hence be approximated to an arbitrary accuracy by sufficiently wide MLPs by virtue of standard universal approximation results [Hornik et al., 1989].
(2) We need to prove that there exist functions $\phi_{1}$ and $\phi_{2}$ such that $\phi_{1}\big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\big{)}\approx SPD^{g}[i,j]$ for every graph $g$ with adjacency matrix $A\in\mathbb{G}_{n}$ and every node-pair $(i,j)$ in $g$.
If we can ensure that we can choose $\phi_{2}$ such that it is possible to map every value $\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})$ to the unique value $SPD^{g}[i,j]$, then we can simply define $\phi_{1}:\mathbb{R}\rightarrow\mathbb{R}$ as the piecewise-linear continuous function:
$\phi_{1}\Big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=\phi_{1}\Big{(}\sum_{k=1}^{n}u^{g}_{k}[i]u^{g}_{k}[j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=SPD^{g}[i,j]$ (13)
@@
In this case, by virtue of 12, $SPD^{g}[i,j]=SPD^{g}[p,q]$ since $A^{k}_{sym}[i,j]=A^{k}_{sym}[p,q]$, which implies that there is a $k$-hop path from $i$ to $j$ iff there is a $k$-hop path from $p$ to $q$. Moreover, in this case ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$.
@@
If $SPD^{g}[i,j]=SPD^{g}[p,q]$, we simply set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=SPD^{g}[i,j]$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=SPD^{g}[p,q]$.
If $SPD^{g}[i,j]\neq SPD^{g}[p,q]$ and ${v^{g}_{ij}}^{T}\phi^{g}_{2}\neq{v^{g}_{pq}}^{T}\phi^{g}_{2}$, we are done. If however, ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$, choose an index $r$ such that $v^{g}_{ij}[r]\neq v^{g}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g}_{ij}}^{T}\phi^{g}_{2}$ becomes unequal to ${v^{g}_{pq}}^{T}\phi^{g}_{2}$, in which case we again can set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=SPD^{g}[i,j]$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=SPD^{g}[p,q]$. We are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 13 for other graphs and node-pairs in $\mathbb{G}_{n}$.

### v3

@@ 3.2 Eigenformer architecture
For a layer $l$ of Eigenformer, input node features $x_{i}^{l}\in\mathbb{R}^{d_{\mathcal{V}}}$ for $i=1,...,|\mathcal{V}|$ and input edge features $e_{ij}\in\mathbb{R}^{d_{\mathcal{E}}}$ for $i,j=1,...,|\mathcal{V}|$, the node feature propagation is performed as follows:
${\alpha^{lh}_{spec}}=\phi_{1}^{lh}\Bigl{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\phi_{2}^{lh}(\lambda_{k})\Bigl{)}$ (7)
${\alpha^{lh}_{feat}}=\psi\Bigl{(}\frac{(W^{lh}_{Q}x_{i}^{l})^{T}(W^{lh}_{K}x_{j}^{l})}{\sqrt{d/H}}\Bigl{)}+W^{lh}_{R}\rho\bigl{(}W^{lh}_{A}e_{ij}\bigl{)}$ (8)
$\alpha^{lh}[i,j]=softmax_{j\in\mathcal{V}}({\alpha^{lh}_{spec}}+{\alpha^{lh}_{feat}})$ (9)
@@
$\alpha_{feat}$ is an optional attention component that we introduce to study the gains that node and edge feature attention brings over solitary spectral attention $\alpha_{spec}$. We do not include edge-feature propagation due to the added complexity and marginal benefits in performance. Notably however, we found degree-scaling followed by batch-normalization to be highly effective in incorporating node-degree information. The final architecture of a layer of Eigenformer is illustrated in Figure 2(b) along with the spectrum-aware attention mechanism that it uses in Figure 2(a).
Table 1: Test performance on five benchmarks from Dwivedi et al. [2020]. Due to limited training resources, we show results for only a single training run while other baselines report mean $\pm$ standard deviation of 4 runs with different random seeds. Color codes denote top first, second, third results.
Model ZINC MNIST CIFAR10 PATTERN CLUSTER
@@
GCN 0.367$\pm$.011 90.705$\pm$.218 55.710$\pm$.381 71.892$\pm$.334 68.498$\pm$.976
GIN 0.526$\pm$.051 96.485$\pm$.252 55.255$\pm$1.527 85.387$\pm$.136 64.716$\pm$1.553
GAT 0.384$\pm$.007 95.535$\pm$.205 64.223$\pm$.455 78.271$\pm$.186 70.587$\pm$.447
GatedGCN 0.282$\pm$.015 97.340$\pm$.143 67.312$\pm$.311 85.568$\pm$.088 73.840$\pm$.326
GatedGCN-LSPE 0.090$\pm$.001 - - - -
PNA 0.188$\pm$.004 97.940$\pm$.12 70.35$\pm$.63 - -
DGN 0.168$\pm$.003 - 72.838$\pm$.417 86.680$\pm$.034 -
GSN 0.101$\pm$.010 - - - -
CIN 0.079$\pm$.006 - - - -
CRaW1 0.085$\pm$.004 97.944$\pm$.050 69.013$\pm$.259 - -
GIN-AK+ 0.080$\pm$.001 - 72.19$\pm$.13 86.850$\pm$.057 -
SAN 0.139$\pm$.006 - - 86.581$\pm$.037 76.691$\pm$.65
Graphormer 0.122$\pm$.006 - - - -
K-Subgraph SAT 0.094$\pm$.008 - - 86.848$\pm$.037 77.856$\pm$.104
EGT 0.108$\pm$.009 98.173$\pm$.087 68.702$\pm$.409 86.821$\pm$.020 79.232$\pm$.348
Graphormer-URPE 0.086$\pm$.007 - - - -
Graphormer-GD 0.081$\pm$.009 - - - -
GPS 0.070$\pm$.004 98.051$\pm$.126 72.298$\pm$.356 86.685$\pm$.059 78.016$\pm$.180
GRIT 0.059$\pm$.002 98.108$\pm$.111 76.468$\pm$.881 87.196$\pm$.076 80.026$\pm$.277
EXPHORMER - 98.550$\pm$.039 74.69$\pm$.125 86.740$\pm$.015 78.070$\pm$.037
Eigenformer 0.077 97.142 67.316 86.680 77.262
Eigenformer (feat.) 0.089 98.362 70.194 86.738 77.456
Table 2: Test performance on a benchmark from the long-range graph benchmarks (LRGB) work [Dwivedi et al., 2022b]. Due to limited training resources, we show results for only a single training run while other baselines report mean $\pm$ standard deviation of 4 runs with different random seeds. Color codes denote top first, second, third results.
@@
Eigenformer 0.6414 0.2599
Eigenformer (feat.) 0.6317 0.2687
@@
(c) CIFAR10
(d) PATTERN
(e) CLUSTER
(f) Peptides-func/struct
Figure 3: (Smoothed) potential $\sigma_{k}$ vs eigenvalue $\lambda_{k}$
@@
4 Experimental Results
We evaluate Eigenformer on five benchmarks from the widely used Benchmarking GNNs work [Dwivedi et al., 2020], viz. ZINC, MNIST, CIFAR10, PATTERN and CLUSTER. These datasets include different graph learning objectives like node classification, graph classification and graph regression. They also cover widely different domains ranging from chemistry and computer vision to synthetically generated graphs. Further, we report numbers on two benchmarks from the Long-Range Graph Benchmark (LRGB) [Dwivedi et al., 2022b] to gauge how well SAA does in capturing long-range dependencies. Further details about the experimental setup can be found in appendix B.
Baselines: We compare our method to a variety of SOTA GNN methods including (hybrid) Graph Transformers: GraphGPS Rampášek et al. [2022], GRIT Ma et al. [2023], EGT Hussain et al. [2022], SAN Kreuzer et al. [2021], Graphormer Ying et al. [2021], K-Subgraph SAT Chen et al. [2022b] and EXPHORMER Shirzad et al. [2023], MP-GNNs: GCN Kipf and Welling [2017], GAT Veličković et al. [2018], GIN Xu et al. [2019], PNA Corso et al. [2020], GatedGCN Bresson and Laurent [2018], DGN Beaini et al. [2021], GSN Bouritsas et al. [2023] and other performant models like CIN Bodnar et al. [2021], CRaW1 Tönshoff et al. [2021] and GIN-AK+ Zhao et al. [2021].
@@ 4 Experimental Results
Benchmarking GNNs [Dwivedi et al., 2020]: Table 1 compares different methods under the parameter budget used in Rampášek et al. [2022] ($\sim$500K parameters for ZINC, PATTERN and CLUSTER, $\sim$100K parameters for MNIST and CIFAR10). Eigenformer routinely beats all the MP-GNNs and performs comparably to popular Graph Transformers that use different PEs like SAN, Subgraph-SAT, Graphormer, GRIT etc. Further, Eigenformer with both spectral and feature attention does marginally better for these datasets except on ZINC, where spectral attention seems to be adequate. Feature attention appears to be most beneficial for the two directed-graph datasets: MNIST and CIFAR10.
Long-range graph benchmark [Dwivedi et al., 2022b]: We also report numbers on two benchmarks from the LRGB work [Dwivedi et al., 2022b] viz. Peptides-func (multilabel graph classification) and Peptides-struct (graph regression) in Table 2. Again, Eigenformer beats all the MP-GNN models comfortably and is competitive with other Graph Transformers albeit falling short of featuring among the best performers. Interestingly, feature attention is detrimental to performance for these long-range interactions. In conjunction with the results from Table 1, our experiments suggest that feature information is not very helpful in determining how messages ought to be exchanged between nodes.
@@
Visualization of node potentials: It is instructive to visualize the difference in the distribution of potentials (as defined in 5) between adjacent and non-adjacent nodes. Figure 3(f) shows the (smoothed, average) potential distribution w.r.t frequency for all the datasets used in our work. In general, adjacent nodes have higher potentials at the lower end of the spectrum that vary inversely with frequency. Non-adjacent nodes on the other hand, have smaller potentials which increase towards the higher end of the spectrum. We posit that these distributional differences allow the attention mechanism to learn to discern local interactions from long-range interactions.
Visualization of learned attention matrices: Figure 4 shows the adjacency matrices (1-hop: 4(a), 5-hop: 4(b)) and the shortest-path distance matrix (4(c)) for an example graph from the ZINC dataset, along with visualizations of learned attention patterns from the trained model that most closely mimic the corresponding neighborhood pattern (4(d), 4(e), 4(f)). For example, Figure 4(e) shows the attention pattern of a head from Layer 11 of the learned model, which seemingly tries to reproduce the 5-hop neighborhood interactions similar to those shown in Figure 4(b), providing empirical evidence that Eigenformer can learn k-hop interactions.
Figure 4: k-hop neighborhood, shortest-path distance and learned attention matrices for an example graph from the ZINC dataset
Limitations: A couple of open challenges remain however, including the final test performance and computational complexity of Eigenformer. With regards to performance, we believe a more extensive hyperparameter search coupled with edge-propagation might aid in closing the gap to the current SOTA in Graph Transformers. The main bottleneck w.r.t complexity is equation 7 which can be implemented using in-place operations with $O(N^{2})$ memory but in $O(N^{3})$ time. The payoff for the added complexity is increased representational capacity, as SAA can learn $k$-hop interactions for any $k\in\mathbb{Z}$ as against upper-bounded $k$-hop neighborhoods in previous works. An obvious way to reduce the time complexity to $O(N^{2}k)$ is by using a subset of the frequencies. We test this intervention on the ZINC dataset, by randomly choosing $k$ out of $N$ eigenvalues for each example in the data at each training step. Table 4 and Figure 5 show the change in final test performance with varying $k$. The test performance remains impressive even when using $\sim 40\%$ of the eigenvalues, suggesting a graceful degradation in performance with decreasing number of eigenvalues ($k$) used.
Table 3: Test MAE for the ZINC dataset for varying number (k) of eigenvalues used. N=37.
k k/N Test MAE $\downarrow$
37 1.0 0.089
30 0.81 0.097
22 0.59 0.108
15 0.41 0.133
8 0.22 0.180
Figure 5: Percentage change in MAE with decreasing number (k) of eigenvalues used
5 Related Work
Considering the importance of graph connectivity/topology information for any GNN architecture, there have been many attempts to incorporate graph inductive biases in Transformers following the introduction of Graph Transformers in Li et al. [2019].
Graph inductive biases in Transformers: Dwivedi and Bresson [2020] performs attention only over the 1-hop neighborhood of nodes in addition to the use of Laplacian eigenvectors as node PEs. With Graph-BERT, Zhang et al. [2020] proposes to train Transformers using sampled linkless (no edges) subgraphs within their local contexts, designed with an emphasis on parallelization and mainly as a pretraining technique. They use a number of PEs like the WL Absolute Role Embedding, Intimacy-based Relative Positional Embedding and Hop-based Relative Distance Embedding, in order to retain the original graph structural information in the linkless subgraphs.
Kreuzer et al. [2021] uses a Transformer encoder to create learnable PEs from a combination of eigenvalues and eigenvectors which are then fed to the main Transformer network. Dwivedi et al. [2022a] decouples the node features and the node PEs with a separate propagation step for the PEs. They consider two different initializations for the PEs: Laplacian PE (LapPE) and Random Walk PE (RWPE), and show that RWPE outperforms LapPE, indicating that learning the sign invariance ambiguity in LapPE is more difficult than not exactly having unique node representation for each node. Ying et al. [2021] proposes the use of a fixed centrality encoding to inject degree information in their Graphormer.
With GraphiT, Mialon et al. [2021] enumerates, encodes and aggregates local sub-structures such as paths of length-k using Graph convolutional kernel networks (GCKN) [Chen et al., 2020] to create node PEs. Ma et al. [2023] proposes GRIT, again using PEs initialized with random walk probabilities but makes them flexible and learnable by composing with an MLP, and provide theoretical proof of the expressivity of such a composition. Further, they inject degree information using an adaptive degree-scaler [Corso et al., 2020] in every layer, noting that attention mechanisms are invariant to node-degrees.
Structure-aware attention mechanisms: More similar to our work, there has also been research on creating structure-aware attention mechanisms in order to infuse node structural and neighborhood similarity information directly in the attention logic. Mialon et al. [2021] views self-attention as kernel smoothing and bias the attention scores using diffusion or k-step random-walk kernels. Chen et al. [2022b] argues that such a modification is deficient in filtering out structurally dissimilar nodes, and consider a more generalized kernel accounting for local substructures like k-subtrees and k-subgraphs, around each node. In a similar approach, G et al. [2023] aggregates node feature information from k-subtrees but augment their Transformer with a global node and train positional encodings using the Neighbourhood Contrastive Loss (NCE). Ying et al. [2021] proposed to inject spatial and edge information by adding learnable encodings to the attention scores before normalization, while [Menegaux et al., 2023] augments that idea further by learning a different attention score per hidden dimension in their so-called Chromatic Self-Attention (CSA). Learned edge encodings are also incorporated in Hussain et al. [2022] both as a bias to the attention score before normalization and as a sigmoid-gate after normalization.
In contrast to previous work, to the best of our knowledge, there have been no attempts to factorize the attention matrix solely into fixed spectral potentials and learnable frequency importances.
@@ 6 Conclusion
Observing the wide variety of positional encoding schemes for Graph Transformers in the literature, we asked the question whether such encodings are essential to ace graph representational learning objectives. We introduced a novel spectrum-aware attention (SAA) mechanism that incorporates structural graph inductive biases while remaining flexible enough to learn from features in the data. SAA was also shown to be invariant to sign or basis invariance issues that plagued previous PE design strategies. We proposed Eigenformer, a Graph Transformer with SAA at its core, that can be used for a variety of graph learning objectives with different task-specific heads. We empirically test the feasibility of our method through experiments on a number of common GNN benchmarks and find that it performs comparably to many popular GNN architectures employing different schemes for PE generation. We hope this work provides the impetus to the broader graph representational learning research community to pursue the development of other GNN architectures without the use of PEs.
Appendix A Proofs
A.1 Proposition 1
@@
$L_{norm}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{norm}$
Since $L_{norm}$ is real and symmetric, it is diagonalizable and admits an eigendecomposition as follows:
$L_{norm}=U\Lambda U^{-1}$
where the columns of $U=[u_{1},u_{2},...,u_{|\mathcal{V}|}]$ are the eigenvectors of $L_{norm}$. It follows that,
$A_{norm}=I-U\Lambda U^{-1}=U(I-\Lambda)U^{-1}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}(1-\lambda_{i})$ (14)
@@
$\sum_{k=0}^{m}\theta_{k}A^{k}_{norm}=\sum_{i=1}^{|\mathcal{V}|}u_{i}u_{i}^{T}\sum_{k=0}^{m}\theta_{k}(1-\lambda_{i})^{k}$ (15)
The continuous functions $\phi_{1}(x)=x$ and $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ in 6 can hence be approximated to an arbitrary accuracy by sufficiently wide MLPs by virtue of standard universal approximation results [Hornik et al., 1989].
(2) We need to prove that there exist functions $\phi_{1}$ and $\phi_{2}$ such that $\phi_{1}\big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\big{)}\approx f(SPD^{g}[i,j])$ for every graph $g$ with adjacency matrix $A\in\mathbb{G}_{n}$ and every node-pair $(i,j)$ in $g$.
If we can ensure that we can choose $\phi_{2}$ such that it is possible to map every value $\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})$ to the unique value $f(SPD^{g}[i,j])$, then we can simply define $\phi_{1}:\mathbb{R}\rightarrow\mathbb{R}$ as the piecewise-linear continuous function:
$\phi_{1}\Big{(}\sum_{k=1}^{n}\sigma^{g}_{k}[i,j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=\phi_{1}\Big{(}\sum_{k=1}^{n}u^{g}_{k}[i]u^{g}_{k}[j]\cdot\phi_{2}(\lambda^{g}_{k})\Big{)}=f(SPD^{g}[i,j])$ (16)
@@
In this case, by virtue of 15, $SPD^{g}[i,j]=SPD^{g}[p,q]$ since $A^{k}_{norm}[i,j]=A^{k}_{norm}[p,q]$, which implies that there is a $k$-hop path from $i$ to $j$ iff there is a $k$-hop path from $p$ to $q$. Moreover, in this case ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$.
@@
If $f(SPD^{g}[i,j])=f(SPD^{g}[p,q])$, we simply set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=f(SPD^{g}[i,j])$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=f(SPD^{g}[p,q])$.
If $f(SPD^{g}[i,j])\neq f(SPD^{g}[p,q])$ and ${v^{g}_{ij}}^{T}\phi^{g}_{2}\neq{v^{g}_{pq}}^{T}\phi^{g}_{2}$, we are done. If however, ${v^{g}_{ij}}^{T}\phi^{g}_{2}={v^{g}_{pq}}^{T}\phi^{g}_{2}$, choose an index $r$ such that $v^{g}_{ij}[r]\neq v^{g}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g}_{ij}}^{T}\phi^{g}_{2}$ becomes unequal to ${v^{g}_{pq}}^{T}\phi^{g}_{2}$, in which case we again can set $\phi_{1}({v^{g}_{ij}}^{T}\phi^{g}_{2})=f(SPD^{g}[i,j])$ and $\phi_{1}({v^{g}_{pq}}^{T}\phi^{g}_{2})=f(SPD^{g}[p,q])$. We are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$ as the set of such constraints is finite and bounded.